OpenStudy (anonymous):

4. Find the value of x such that the data set has the given mean. 99, 86, 76, 95, x; mean 91 (1 point) 5. Find the value of x such that the data set has the given mean. 31.7, 42.8, 26.4, x; mean 35 (1 point) 6. Find the mean, median, mode, and range of the data set a쪼er you perform the given operation on each data value. 9, 7, 12, 13, 9, 3; add 5 (1 point) mean: 8.8, median: 9, mode: 9, range: 10 mean; 8.8, median: 9, mode: 9, range: 5 mean: 13.8, median: 14, mode: 14, range: 5 mean: 13.8, median: 14, mode: 14, range: 10

OpenStudy (faiqraees):

Sum of all numbers/ frequency = mean (99+ 86+76+ 95+ x)/5 = 91 x = ???

OpenStudy (anonymous):

4. A) x.91 B)x.87 c)x.82 d)x.99

OpenStudy (welshfella):

for the second question solve the equation 4* 35 = 31.7 + 42.8 + 26.4 + x

OpenStudy (anonymous):

maddi i hate school too

OpenStudy (anonymous):

5. A)x=39.1 B)x=42.6 C)x=40.5 D)x= 36.8

OpenStudy (anonymous):

for the first one i know the answer its D

OpenStudy (anonymous):

6. Find the mean, median, mode, and range of the data set aĀer you perform the given operation on each data value. 9, 7, 12, 13, 9, 3; add 5 (1 point) mean: 8.8, median: 9, mode: 9, range: 10 mean; 8.8, median: 9, mode: 9, range: 5 mean: 13.8, median: 14, mode: 14, range: 5 mean: 13.8, median: 14, mode: 14, range: 10

OpenStudy (welshfella):

to get the median first arrange the numbers in ascending order the median is the middle number in the list.

OpenStudy (anonymous):

i dont get the add 5, is it just like +5 at the end thats whats confusing me

OpenStudy (welshfella):

yes just add 5 to the list so we gat 9 7 12 9 13 3 5 7 numbers

OpenStudy (welshfella):

in order thay are 3, 5 - can you continue?

OpenStudy (anonymous):

when i did it without the 5 in there i got 8.8 and when i added the 5 i got 8.2

OpenStudy (welshfella):

whats that the mean ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for the median i got 9

OpenStudy (welshfella):

yes i'm confused too if you add 5 the mean is 8.3 and if you dont its 8.8 as you said

OpenStudy (anonymous):

yes

OpenStudy (welshfella):

the median is 9 whether you add the 5 or not lol!

OpenStudy (welshfella):

maybe you have to add 5 to each of the values

OpenStudy (welshfella):

that would make the list 8,12,14,14,17,18 yea thats what it is

OpenStudy (welshfella):

the median of this is 14 . also the mode ( the most frequent occurring number) is also 14

OpenStudy (welshfella):

can you work out the mean and the range?

OpenStudy (anonymous):

i thnk so

OpenStudy (welshfella):

the range is highest number - lowest number

OpenStudy (anonymous):

so 10

OpenStudy (welshfella):

yea

OpenStudy (anonymous):

so 10

OpenStudy (anonymous):

so the answer is D

OpenStudy (welshfella):

right the mean is 13.8

OpenStudy (welshfella):

D

OpenStudy (anonymous):

Thanks so much!

OpenStudy (welshfella):

yw